OpenStudy (anonymous):

I am having trouble with this please help Points A and B lie on the graph of the equation y = 7x. What are the y-coordinates for each point? Point A: (−2, ) Point B: (4, )

jimthompson5910 (jim_thompson5910):

point A has an x coordinate of -2 x = -2 replace the x in the equation with -2 and evaluate y = 7*x y = 7*(-2) y = ???

OpenStudy (anonymous):

-14

jimthompson5910 (jim_thompson5910):

so point A is (-2, -14)

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

the y coordinate of point A is -14

jimthompson5910 (jim_thompson5910):

you'll do the same for point B

OpenStudy (anonymous):

28

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

And I have more

jimthompson5910 (jim_thompson5910):

ok post as a new question. It's a good idea to keep one problem per post

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

please

jimthompson5910 (jim_thompson5910):

you can tag me in the new question when you make a new post